[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Jun 16, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Fri Jun 16 14:58:47 UTC 2017


Here is today's summary of NHC computer operations:
--- Matt Sardi

----------

TAFB:
1. The primary compute server (nhc-lx-comp01) was rebooted by NCO  
yesterday, which resolved the Gempak 'allocate color' errors that were  
causing the Danger graphics to fail. The Danger graphics scripts are  
running on comp01 again, I switched them over to nhc-lx-comp03 on Wednesday  
as a temporary work-around.

2. AWIPS workstation LX5 screen was blank upon login. A reboot was  
performed, which resolved the problem.

3. Cumulus was frozen. A hard reboot was performed.
HSU:
1. GOES-16 CONUS VIS red band has stripes of missing data. It was  
particularly bad around 14-15 UTC. Monica Bozeman indicates that the 14-15  
UTC disruption was due to a NESDIS calibration/validation test.

2. Specialists continue to report performance issues on the Atlantic  
N-AWIPS workstation RIDGES. Dave Zelinsky opened a helpdesk ticket to  
change the 'color depth' from 24 back to 8 consistent with the other  
workstations.
Other items:
1. On Monday, GOES-16 processing will be switched over to our secondary  
GRB02 server temporarily while the primary GRB01 server is upgraded with  
new Python code. Please let Monica know if there are data issues (missing  
data, etc.) when this switchover occurs.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170616/a2eed7af/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list